OpenStudy (anonymous):

Tell whether the lines for each pair of equations are parallel, perpendicular, or neither. 3x - 5y = 3 -5x + 3y = 8

OpenStudy (anonymous):

slope of \(ax+by=c\) is \(-\frac{a}{b}\) compute both of these. you will see that one is the negative reciprocal of the other.

OpenStudy (anonymous):

so is it perpendicular

OpenStudy (anonymous):

so i had it right @Chlorophyll lol

OpenStudy (lgbasallote):

negative reciprocal? i thinkim seeing just reciprocal lol

OpenStudy (anonymous):

Tell whether the lines for each pair of equations are parallel, perpendicular, or neither. 3x - 5y = 3 -5x + 3y = 8 a. parallel b. perpendicular c. neither

OpenStudy (lgbasallote):

first equation: -a/b = -3/-5 = 3/5 second equation: -a/b = -(-5)/3 = 5/3

OpenStudy (lgbasallote):

so what do you think?

OpenStudy (anonymous):

Hint: they're NOT: equal they're NOT reciprocal ( m1 * m2 ≠ -1)

OpenStudy (anonymous):

@davonnn Common, you can do it :)

OpenStudy (anonymous):

so neither ?

OpenStudy (lgbasallote):

yep..it's neither yayy

OpenStudy (anonymous):

Why neither including "?" :'(